You are standing on the roadway of the Ambassador Bridge. It is 46.3 m above the water below. You throw a rock straight out with a velocity of 10.7 m/s. What is the horizontal distance that the rock will travel before it hits the water?
1
Expert's answer
2012-11-05T09:57:46-0500
At first we calculate time itneeds to fall: t=sqrt(2h/g)=3.07s Horizontal distance it covers in this time isS=tv=3.07*10.7=32.87m.